# Analog model of Kerr black holes in scalar-tensor theories using binary Bose-Einstein condensates: Quasibound states and tachyonic instabilities

### v1

@@ II FORMALISM
as $r_{0}\gg r_{H}$ with $\zeta=r_{0}-\alpha^{-1}{\ln\left(\frac{2\sqrt{2d\Omega_{0}}}{\omega_{R}}\right)}$ and $\Omega_{0}>0$, from which $\omega\simeq\omega_{R}$ can be determined. In the limit of a thin width of the mass shell where $\alpha\rightarrow\infty$, $\zeta\rightarrow r_{0}$, the real part of the frequency becomes $\omega_{R}\sim(\frac{m}{2}+n+\frac{1}{2})\pi/r_{0}$ reducing to the known result in [50]. After plugging in the resulting frequency $\omega\simeq\omega_{R}$ in $V_{\text{eff}}$, increasing $r_{0}$ shifts the barrier away from the rotational barrier so that the shaded area in Fig. 2 decreases, giving a smaller value of $\omega_{R}$.
@@
Nevertheless, it is noted that, for example, for $n=0$ (the fundamental mode) and $n=1$ (the overtone mode), if $r_{0}$ is above a certain value shown in Fig. 2, the imaginary part of the eigenfrequency $\omega_{I}$ becomes positive, indicating superradiant instabilities due to the existence of the so-called quasibound state. Then, as $r_{0}$ falls below the critical value, the negativeness of $\omega_{I}$ reveals the quasinormal mode. On the contrary, as $\Omega_{0}$ increases, the shaded area increases, resulting in a larger value of $\omega_{R}$. Similarly, there is a critical value of $\Omega_{0}$, above (below) which the mode suffers (becomes) superradiant instabilities (quasinormal modes). In this setup, sending a monochromatic wave into the vortex will be scattered even if the mass shell shields the vortex. If its frequency is $\omega<m\ell$, it will cause superradiant amplification, and if the frequency is also close to the spectrum frequency $\omega\sim\omega_{R}$, significant resonance amplification will occur, as shown in Fig. 2. Either a stable quasinormal mode or an unstable quasibound state is partially trapped within the mass shell and partially propagate towards spatial infinity, resembling a new class of modes in the astrophysics [26]. The instabilities associated with the quasibound states are also evident from the growth of the radial profile with time in Fig. 2. While the function $H(r=2,t)$ grows, the echo time of two neighboring peaks is about $2r_{0}$. Within the lifetime of the condensates, instabilities will destabilize the background solutions, where backreaction effects must be systematically taken into account.
@@
Figure 3: The left panel shows the comparison among numerics, WKB and PT methods for the imaginary part of frequency as a function of $\Omega_{0}$. The right top panels depict the effective potential and the normalized radial profile for modes $n=0,1$. The right bottom panel reveals the time-dependent profile with varying $\Omega_{0}$. The parameters are fixed as $m=1,\,r_{0}=10,$ and $\alpha=1/2$.
@@
Tachyonic instabilities. A negative mass squared could trigger tachyonic instabilities in scalar-tensor theories, which further leads to spontaneous scalarization. Tachyonic instabilities arise from a completely different mechanism than that of superradiant instabilities. It is expected that tachyonic instabilities lead to a large imaginary part with a short lifetime, namely $\omega_{I}\gg\omega_{R}$. We present numerics, PT, and WKB analytical methods to confirm the existence of the tachyonic instability spectrum in Fig. 3. A fundamental mode ($n=0$) has the largest frequency $\omega\simeq\omega_{I}$ for $\omega_{I}>0$ compared with the overtones. A larger coupling $\Omega_{0}$ and a wider shell $\alpha^{-1}$ lead to stronger tachyonic instabilities. The radial profiles $H_{m}(r,\omega)$ are also drawn for $n=0$, the fundamental mode, and $n=1$, the overtone mode, which lie within the respective classical allowed regimes. Let us assume that (7) can still be applied to give an analytical formula for the eigenfrequency with $\omega\simeq i\omega_{I}$. Substituting $\omega\simeq i\omega_{I}$ into $V_{\text{eff}}$ shows the potential profile in Fig. 3. The shaded area for the classical allowed region obeys (7) for an integer $n$ given by
@@
Conclusion We conclude this letter by summarizing the parameter constraints for a successful analog gravity model. The zero-phase difference between the two types of condensates give perturbed fields with positive mass squared, and the $\pi$-phase difference can give negative mass squared where the choice of one of two modes allows for experimental feasibility [32, 33]. Positive sound speed squared is required, where $g_{12}/g<1\pm\Omega_{0}/\rho g$ (with a dimension-restored variable $\Omega_{0}$), leading to a further constraint on $\Omega_{0}<\rho(g-g_{12})$ in the $\pi$-phase difference modes [51, 52]. The specific form of the mass shell function in (6) that we proposed here illustrates the analog model. The same instabilities can be observed for the general parametrization of the mass distribution by its typical distance from the black hole’s horizon and the width of the distribution. The experimentally challenging but achievable spatially dependent Rabi coupling strength has been discussed in [36].
Acknowledgements.
@@
SUPPLEMENTAL MATERIAL
Background solutions and Thomas-Fermi approximations
We show the detailed derivations of the background solutions for the density and phase of two-component BECs. Plugging in the parametrization of the condensate wavefunction of $\langle\hat{\Psi}_{i}\rangle(\mathbf{r},t)=\sqrt{\rho_{i}(\mathbf{r},t)}\,e^{i\theta_{i}(\mathbf{r},t)-i\mu_{i}t}$ to (1), the density and phase satisfy the equations given respectively by
$i\partial_{t}\left(\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}\right)=\left(-\frac{1}{2m_{a}}\nabla^{2}+V_{\text{ext},1}+g_{11}\rho_{1}+g_{12}\rho_{2}\right)\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}-\frac{\Omega}{2}\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)},$ (S1)
$i\partial_{t}\left(\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)}\right)=\left(-\frac{1}{2m_{a}}\nabla^{2}+V_{\text{ext},2}+g_{22}\rho_{2}+g_{12}\rho_{1}\right)\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)}-\frac{\Omega}{2}\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}.$ (S2)
@@
$g_{11}=g_{22}\equiv g,\quad g_{12}<g,$ (S3)
$\mu_{1}=\mu_{2}\equiv\mu,\quad V_{1}=V_{2}\equiv V_{0}.$ (S4)
@@
$i\partial_{t}\left(\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}\right)=\left(-\frac{1}{2m_{a}}\nabla^{2}+V_{0}+g\rho_{1}+g_{12}\rho_{2}\right)\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}-\frac{\Omega}{2}\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)},$ (S5)
the left-hand side of (S5) gives
$i\left[\partial_{t}\sqrt{\rho_{1}}+i\sqrt{\rho_{1}}(\partial_{t}\theta_{1}-\mu)\right]e^{i(\theta_{1}-\mu t)}$ (S6)
and the right-hand side of (S5) becomes
@@
$\quad\qquad V_{0}+g_{11}\rho_{1}+g_{12}\rho_{2}\Big{\}}\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu t)}-\frac{1}{2}\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu t)}.$ (S7)
Therefore, Its imaginary part of (S5) leads to
@@
$\dot{\rho}_{1}=-\frac{1}{m_{a}}\Big{(}2\nabla\sqrt{\rho_{1}}\cdot\nabla\theta_{1}+\sqrt{\rho_{1}}\nabla^{2}\theta_{1}\Big{)}\sqrt{\rho_{1}}-{\Omega}\sqrt{\rho_{1}\rho_{2}}\sin{(\theta_{2}-\theta_{1})}\,,$ (S8a)
@@
$\dot{\theta}_{1}=\frac{1}{2m_{a}\sqrt{\rho_{1}}}\left[\nabla^{2}\sqrt{\rho_{1}}-\sqrt{\rho_{1}}(\nabla\theta_{1})^{2}\right]-\left(V_{0}+g\rho_{1}+g_{12}\rho_{2}-\mu\right)+\frac{\Omega}{2}\sqrt{\frac{\rho_{2}}{\rho_{1}}}\cos{(\theta_{2}-\theta_{1})}.$ ) . (S8b)
@@
$\dot{\rho}_{2}=-\frac{1}{m_{a}}\Big{(}2\nabla\sqrt{\rho_{2}}\cdot\nabla\theta_{2}+\sqrt{\rho_{2}}\nabla^{2}\theta_{2}\Big{)}\sqrt{\rho_{2}}-{\Omega}\sqrt{\rho_{1}\rho_{2}}\sin{(\theta_{1}-\theta_{2})},$ (S9a)
$\dot{\theta}_{2}=\frac{1}{2m\sqrt{\rho_{2}}}\left[\nabla^{2}\sqrt{\rho_{2}}-\sqrt{\rho_{2}}(\nabla\theta_{2})^{2}\right]-\left(V_{0}+g\rho_{2}+g_{12}\rho_{1}-\mu\right)+\frac{\Omega}{2}\sqrt{\frac{\rho_{1}}{\rho_{2}}}\cos{(\theta_{1}-\theta_{2})}.$ ) . (S9b)
@@
$\rho_{1}=\rho_{2}=\rho,\qquad\theta_{1}=\theta\,,\theta_{2}=\theta\,;\quad\quad\theta_{1}=\theta\,,\theta_{2}=\theta\pm\pi\,$ (S10)
for zero-phase difference and $\pi$-phase difference. Imposing (S10) into (Background solutions and Thomas-Fermi approximations) and (Background solutions and Thomas-Fermi approximations) gives
@@
$\qquad\qquad\Rightarrow-\nabla\cdot\left(\rho\frac{\nabla\theta}{m_{a}}\right)=0\Rightarrow\nabla\cdot(\rho\vec{v})=0,$ (S11a)
@@
$\qquad\qquad\Rightarrow\frac{1}{2m_{a}\sqrt{\rho}}\nabla^{2}\sqrt{\rho}-\frac{m_{a}}{2}\mathbf{v}^{2}-V_{0}-(g+g_{12})\rho+\mu\pm\frac{\Omega}{2}=0,$ (S11b)
where we have used the definition of the condensate flow velocity $\mathbf{v}=\gradient\theta/m_{a}$. The first term in (S11b) is the quantum pressure to be ignored in the hydrodynamical regime, and the plus and minus signs of the last term correspond to the solutions of zero and $\pi$ phase differences, respectively.
@@
$\mathbf{v}=-\frac{d}{m_{a}r}\mathbf{e}_{r}+\frac{\ell}{m_{a}r}\mathbf{e}_{\phi},$ (S12)
then in the hydrodynamical limit (S11b) becomes
$-\left(\frac{d^{2}+\ell^{2}}{2m_{a}}\right)\frac{1}{r^{2}}-V_{0}-(g+g_{12})\rho+\mu\pm\frac{\Omega}{2}=0.$ (S13)
@@
$\rho(r)=\rho_{\infty}(1-\frac{R_{\text{TF}}^{2}}{r^{2}})$ (S14)
@@
$\rho_{\infty}=\frac{1}{g+g_{12}}\left(\mu-V_{0}\pm\frac{\Omega}{2}\right),$ (S15)
@@
$R_{\text{TF}}^{2}=\frac{d^{2}+\ell^{2}}{2m_{a}\rho_{\infty}(g+g_{12})}.$ (S16)
Pöschl-Teller method and Continued fraction method

### v2

@@ II FORMALISM
as $r_{0}\gg r_{H}$ with $\zeta=r_{0}-\alpha^{-1}{\ln\left(\frac{2\sqrt{2d\Omega_{0}}}{\omega_{R}}\right)}$ and $\Omega_{0}>0$, from which $\omega\simeq\omega_{R}$ can be determined. In the limit of a thin width of the mass shell where $\alpha\rightarrow\infty$, $\zeta\rightarrow r_{0}$, the real part of the frequency becomes $\omega_{R}\sim(\frac{m}{2}+n+\frac{1}{2})\pi/r_{0}$ reducing to the known result in Cardoso et al. (2004b). After plugging in the resulting frequency $\omega\simeq\omega_{R}$ in $V_{\text{eff}}$, increasing $r_{0}$ shifts the barrier away from the rotational barrier so that the shaded area in Fig. 2 decreases, giving a smaller value of $\omega_{R}$.
@@
Nevertheless, it is noted that, for example, for $n=0$ (the fundamental mode) and $n=1$ (the overtone mode), if $r_{0}$ is above a certain value shown in Fig. 2, the imaginary part of the eigenfrequency $\omega_{I}$ becomes positive, indicating superradiant instabilities due to the existence of the so-called quasibound state. Then, as $r_{0}$ falls below the critical value, the negativeness of $\omega_{I}$ reveals the quasinormal mode. On the contrary, as $\Omega_{0}$ increases, the shaded area increases, resulting in a larger value of $\omega_{R}$. Similarly, there is a critical value of $\Omega_{0}$, above (below) which the mode suffers (becomes) superradiant instabilities (quasinormal modes). In this setup, sending a monochromatic wave into the vortex will be scattered even if the mass shell shields the vortex. If its frequency is $\omega<m\ell$, then it will cause superradiant amplification, and if the frequency is also close to the spectrum frequency $\omega\sim\omega_{R}$, significant resonance amplification will occur, as shown in Fig. 2. Either a stable quasinormal mode or an unstable quasibound state is partially trapped within the mass shell and partially propagate towards spatial infinity, resembling a new class of modes in the astrophysics Cardoso et al. (2013b). The instabilities associated with the quasibound states are also evident from the growth of the radial profile with time in Fig. 2. While the function $H_{m}(r=2,t)$ grows, the echo time of two neighboring peaks is about $2r_{0}$. Within the lifetime of the condensates, instabilities will destabilize the background solutions, where backreaction effects must be systematically taken into account.
Figure 3: The absolute square of the scattering amplitude as a function of $\tilde{\phi}$ for frequency $\omega=0.5$, with $\Omega_{0}=0.2$ (red)(quasibound state frequency shown in Fig. 2), and $\Omega_{0}=0.1$ (blue), $\Omega_{0}=0.0$ (green). The interference pattern given by the real part of $e^{i\omega x}+f_{\omega}(\tilde{\phi})\frac{e^{i\omega r^{\ast}}}{\sqrt{r}}$ is shown in inset figures.
Another method to justify the existence of the quasibound states is through scattering experiment where time dependent measurements might face challenges due to its long time scale $\tau\propto 1/|\omega_{I}|$. Therefore, we investigate scattering to see the effect of mass shell around the black hole on the imprint carried by the scattered wave. Scattering problems for the scalar field in the draining bathtub are intensively studied Oliveira et al. (2010); Dolan et al. (2012b); Dolan and Oliveira (2013). The differential cross section is given by
$\frac{d\sigma}{d\tilde{\phi}}=|f_{\omega}(\tilde{\phi})|^{2},$ (12)
where the scattering amplitude $f_{\omega}$ obeys the solution of the Klein-Gordon equation in (2) with the metric (4)
$\delta\theta_{s}(t,r,\tilde{\phi})=e^{-i\omega t}\left[e^{i\omega x}+f_{\omega}(\tilde{\phi})\frac{e^{i\omega r^{\ast}}}{\sqrt{r}}\right].$ (13)
We present the results of the scattering cross section $d\sigma/d\tilde{\phi}$ obtained from the partial wave summation in Fig. 3. We find that the large angular scattering occurs for the scalar field with nonzero mass term compared to the zero mass term cases Dolan et al. (2012b); Dolan (2013). In the case of the quasibound state, the absolute square of the scattered amplitude due to resonance amplification becomes even larger than nonresonance cases. The corresponding the interference patterns that might be observed in experiments given by the real part of $e^{i\omega x}+f_{\omega}(\tilde{\phi})\frac{e^{i\omega r^{\ast}}}{\sqrt{r}}$ are also displayed in the inset figures.
@@
Figure 4: The left panel shows the comparison among numerics, WKB and PT methods for the imaginary part of frequency as a function of $\Omega_{0}$. The right top panels depict the effective potential and the normalized radial profile for modes $n=0,1$. The right bottom panel reveals the time-dependent profile with varying $\Omega_{0}$. The parameters are fixed as $m=1,\,r_{0}=10$ and $\alpha=1/2$.
IV Tachyonic instabilities
@@
A negative mass squared could trigger tachyonic instabilities in scalar-tensor theories, which further leads to spontaneous scalarization. Tachyonic instabilities arise from a completely different mechanism than that of superradiant instabilities. It is expected that tachyonic instabilities lead to a large imaginary part with a short lifetime, namely $\omega_{I}\gg\omega_{R}$. We present numerics, PT, and WKB analytical methods to confirm the existence of the tachyonic instability spectrum in Fig. 4. A fundamental mode ($n=0$) has the largest frequency $\omega\simeq\omega_{I}$ for $\omega_{I}>0$ compared with the overtones. A larger coupling $\Omega_{0}$ and a wider shell $\alpha^{-1}$ lead to stronger tachyonic instabilities. The radial profiles $H_{\omega,m}(r)$ are also drawn for $n=0$, the fundamental mode, and $n=1$, the overtone mode, which lie within the respective classical allowed regimes. Let us assume that (10) can still be applied to give an analytical formula for the eigenfrequency with $\omega\simeq i\omega_{I}$. Substituting $\omega\simeq i\omega_{I}$ into $V_{\text{eff}}$ shows the potential profile in Fig. 4. The shaded area for the classical allowed region obeys (10) for an integer $n$ given by
@@
V Conclusion
We conclude this article by summarizing the parameter constraints for a successful analog gravity model. The zero-phase difference between the two types of condensates give perturbed fields with positive mass squared, and the $\pi$-phase difference can give negative mass squared where the choice of one of two modes allows for experimental feasibility Zibold et al. (2010); Abbarchi et al. (2013). Positive sound speed squared is required, where $g_{12}/g<1\pm\Omega_{0}/\rho g$ (with a dimension-restored variable $\Omega_{0}$), leading to a further constraint on $\Omega_{0}<\rho(g-g_{12})$ in the $\pi$-phase difference modes Bernier et al. (2014); Recati and Stringari (2022). The experimentally challenging but achievable spatially dependent Rabi coupling strength has been discussed in Nicklas et al. (2011).
VI acknowledgments
@@ V Conclusion
Appendix A BACKGROUND SOLUTIONS AND THOMAS-FERMI APPROXIMATIONS
We show the detailed derivations of the background solutions for the density and phase of two-component BECs. Plugging in the parametrization of the condensate wave functions of $\langle\hat{\Psi}_{i}\rangle(\mathbf{r},t)=\sqrt{\rho_{i}(\mathbf{r},t)}\,e^{i\theta_{i}(\mathbf{r},t)-i\mu_{i}t}$ to (1), the density and phase satisfy the equations given, respectively by
$i\partial_{t}\left(\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}\right)=\left(-\frac{1}{2m_{a}}\nabla^{2}+V_{\text{ext},1}+g_{11}\rho_{1}+g_{12}\rho_{2}\right)\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}-\frac{\Omega}{2}\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)},$ (15a)
$i\partial_{t}\left(\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)}\right)=\left(-\frac{1}{2m_{a}}\nabla^{2}+V_{\text{ext},2}+g_{22}\rho_{2}+g_{12}\rho_{1}\right)\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)}-\frac{\Omega}{2}\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}.$ (15b)
@@ Appendix A BACKGROUND SOLUTIONS AND THOMAS-FERMI APPROXIMATIONS
$g_{11}=g_{22}\equiv g,\quad g_{12}<g,$ (16)
$\mu_{1}=\mu_{2}\equiv\mu,\quad V_{1}=V_{2}\equiv V_{0}.$ (17)
@@
$i\partial_{t}\left(\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}\right)=\left(-\frac{1}{2m_{a}}\nabla^{2}+V_{0}+g\rho_{1}+g_{12}\rho_{2}\right)\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu_{1}t)}-\frac{\Omega}{2}\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu_{2}t)},$ (18)
the left-hand side of (18) gives
$i\left[\partial_{t}\sqrt{\rho_{1}}+i\sqrt{\rho_{1}}(\partial_{t}\theta_{1}-\mu)\right]e^{i(\theta_{1}-\mu t)}$ (19)
and the right-hand side of (18) becomes
@@
$\quad\qquad V_{0}+g_{11}\rho_{1}+g_{12}\rho_{2}\Big{\}}\sqrt{\rho_{1}}e^{i(\theta_{1}-\mu t)}-\frac{1}{2}\sqrt{\rho_{2}}e^{i(\theta_{2}-\mu t)}.$ (20)
Therefore, its imaginary part of (18) leads to
@@
$\dot{\rho}_{1}=-\frac{1}{m_{a}}\Big{(}2\nabla\sqrt{\rho_{1}}\cdot\nabla\theta_{1}+\sqrt{\rho_{1}}\nabla^{2}\theta_{1}\Big{)}\sqrt{\rho_{1}}-{\Omega}\sqrt{\rho_{1}\rho_{2}}\sin{(\theta_{2}-\theta_{1})}\,,$ (21a)
@@
$\dot{\theta}_{1}=\frac{1}{2m_{a}\sqrt{\rho_{1}}}\left[\nabla^{2}\sqrt{\rho_{1}}-\sqrt{\rho_{1}}(\nabla\theta_{1})^{2}\right]-\left(V_{0}+g\rho_{1}+g_{12}\rho_{2}-\mu\right)+\frac{\Omega}{2}\sqrt{\frac{\rho_{2}}{\rho_{1}}}\cos{(\theta_{2}-\theta_{1})}.$ ) . (21b)
@@
$\dot{\rho}_{2}=-\frac{1}{m_{a}}\Big{(}2\nabla\sqrt{\rho_{2}}\cdot\nabla\theta_{2}+\sqrt{\rho_{2}}\nabla^{2}\theta_{2}\Big{)}\sqrt{\rho_{2}}-{\Omega}\sqrt{\rho_{1}\rho_{2}}\sin{(\theta_{1}-\theta_{2})},$ (22a)
$\dot{\theta}_{2}=\frac{1}{2m\sqrt{\rho_{2}}}\left[\nabla^{2}\sqrt{\rho_{2}}-\sqrt{\rho_{2}}(\nabla\theta_{2})^{2}\right]-\left(V_{0}+g\rho_{2}+g_{12}\rho_{1}-\mu\right)+\frac{\Omega}{2}\sqrt{\frac{\rho_{1}}{\rho_{2}}}\cos{(\theta_{1}-\theta_{2})}.$ ) . (22b)
@@
$\rho_{1}=\rho_{2}=\rho,\qquad\theta_{1}=\theta\,,\theta_{2}=\theta\,;\quad\quad\theta_{1}=\theta\,,\theta_{2}=\theta\pm\pi\,$ (23)
for zero-phase difference and $\pi$-phase difference. Imposing (23) into (21) and (22) gives
@@
$\qquad\qquad\Rightarrow-\nabla\cdot\left(\rho\frac{\nabla\theta}{m_{a}}\right)=0\Rightarrow\nabla\cdot(\rho\vec{v})=0,$ (24a)
@@
$\qquad\qquad\Rightarrow\frac{1}{2m_{a}\sqrt{\rho}}\nabla^{2}\sqrt{\rho}-\frac{m_{a}}{2}\mathbf{v}^{2}-V_{0}-(g+g_{12})\rho+\mu\pm\frac{\Omega}{2}=0,$ (24b)
where we have used the definition of the condensate flow velocity $\mathbf{v}=\gradient\theta/m_{a}$. The first term in (24b) is the quantum pressure to be ignored in the hydrodynamical regime, and the plus and minus signs of the last term correspond to the solutions of zero and $\pi$ phase differences, respectively.
@@
$\mathbf{v}=-\frac{d}{m_{a}r}\mathbf{e}_{r}+\frac{\ell}{m_{a}r}\mathbf{e}_{\phi},$ (25)
then in the hydrodynamical limit (24b) becomes
$-\left(\frac{d^{2}+\ell^{2}}{2m_{a}}\right)\frac{1}{r^{2}}-V_{0}-(g+g_{12})\rho+\mu\pm\frac{\Omega}{2}=0.$ (26)
@@
$\rho(r)=\rho_{\infty}(1-\frac{R_{\text{TF}}^{2}}{r^{2}})$ (27)
@@
$\rho_{\infty}=\frac{1}{g+g_{12}}\left(\mu-V_{0}\pm\frac{\Omega}{2}\right),$ (28)
@@
$R_{\text{TF}}^{2}=\frac{d^{2}+\ell^{2}}{2m_{a}\rho_{\infty}(g+g_{12})}.$ (29)
In the text, we choose the experimental parameters so that the Thomas-Fermi radius $R_{\text{TF}}$ is far smaller than the horizon radius $r_{H}$, namely $r_{H}\gg R_{\text{TF}}$, where the density can be safely treated as a constant $\rho(r>r_{H})\simeq\rho_{\infty}$ Patrick et al. (2022). Regarding to the density beyond constant approximation, the superradiant amplification could be modified Oliveira et al. (2018); Demirkaya et al. (2020); Cardoso and Richartz (2022).
Appendix B PÖSCHL-TELLER METHOD AND CONTINUED FRACTION METHOD